OpenStudy (anonymous):

add: (3x^2-4xy+y^2)+-8x^2-9xy-y^2)+(x^2+xy-2y^2)

OpenStudy (anonymous):

Sry the 8x^2 should have a parenthesis around it also

OpenStudy (amistre64):

line them up its easier: 3x^2 -4xy +1y^2 -8x^2 -9xy -1y^2 1x^2 +1xy -2y^2 ----------------- -4x^2 -12xy -2y^2

OpenStudy (anonymous):

what happened to the x^2 and xy where did they go, I'm not understanding that part

OpenStudy (amistre64):

-4x^2 -12xy -2y^2 ^^^^ ^^^^ i dont quite understand your response

OpenStudy (anonymous):

in the last part of the equation where did the x go

OpenStudy (anonymous):

i thought you where supposed to do what was in each bracket and then add those together

OpenStudy (amistre64):

\(3x^2\) \( -4xy\) \( +1y^2 \) \(-8x^2\) \( -9xy\) \( -1y^2 \) \(1x^2\) \( +1xy\) \( -2y^2 \) ----------------- \(-4x^2\) \( -12xy \) \(-2y^2\)

OpenStudy (amistre64):

the brackets are as used up as they can get; there is nothing to do inside of them

OpenStudy (anonymous):

ok i put this question in last night and got the answer you gave me but they said the 4x^2 was positive not negative.

OpenStudy (amistre64):

3+1-8 = -4 4-8 = -4 -4 = -4

OpenStudy (anonymous):

k

OpenStudy (amistre64):

this was your question from last night :) same answer http://openstudy.com/users/amistre64#/users/amistre64/updates/4de0628dc9a28b0b14db66c2